# Rigorous Function Calculi in Ariadne

### v2

@@ Abstract
Keywords: Rigorous numerics; Computable analysis; Function calculus; Mathematical software.
1 Introduction
@@
The Ariadne software package (Ariadne, 2002-, ) provides analysis and verification tools for dynamic systems with continuous state spaces. This includes nonlinear hybrid systems in which continuous evolution is interspersed with discrete events triggered by conditions on the continuous state. This is a very general and complex problem, and requires functionality for many different operations, including solution of algebraic and differential equations and of constrained feasibility problems, and requires general-purpose software. The computational kernel of Ariadne was subsequently developed, based on rigorous numerical methods for working with real numbers, functions and sets in Euclidean space. Previous publications (Benvenuti et al., 2008; Collins et al., 2012) have focused on the use of Ariadne for reachability analysis of systems. The purpose of this paper is to describe the low-level functionality of Ariadne that has been used to build up the high-level algorithms, with a focus on the support for working with Euclidean functions.
@@
1.1 Other Tools and Approaches
@@
  • Ariadne, 2002-  (Collins, Geretti, Villa et al.) for verification of hybrid systems (C++ & Python).
@@
  • iRRAM, 2000-  (Müller, Brauße) for real number arithmetic (C++).
@@
  • AERN, 2005-  (Konec̆ný et al.) for effective real computation (Haskell).
@@
  • ERC (Ziegler, Park et al. Brauße et al. (2018)): a language for exact real computation.
The iRRAM package started off as a tool for efficient and straightforward, high-precision computation with real numbers (Müller, 2001), but more recent developments have extended the functionality to include a function calculus (Brauße et al., 2016) and solution of differential equations. The AERN package is similar in scope to Ariadne, including a function calculus Duracz et al. (2014). A motivation for the development of ERC as a computer analysis (as opposed to algebra) system was given in Brauße et al. (2022).
@@
  • COSY Infinity (Makino and Berz, 2006), the seminal package in which Taylor models were developed Makino and Berz (2003).
  • CAPD Library (Kapela et al., 2021), a comprehensive library for analysis of dynamical systems.
  • Flow* (Chen et al., 2013), a tool for hybrid system evolution which also has Taylor models with interval coefficients.
  • JuliaReach (Bogomolov et al., 2019), a tool for set-based reachability analysis.
  • DynIBEX (Alexandre dit Sandretto and Chapoutot, 2016), has a custom class for trajectories.
  • CORA (Althoff et al., 2018) has Taylor models, but operations use floating-point approximations.
@@
  • Ellipsoidal calculus (Kurzhanski and Varaiya, 2002)
  • Set-valued viability (Cardaliaguet et al., 1999)
  • Taylor-ellipsoid models (Houska et al., 2013)
Since the main aim of this paper is to introduce the framework and capabilities of Ariadne, we do not give a full comparison with these other tools here. For a comparison of the performance of Ariadne against other tools on benchmark problems, see (Geretti et al., 2020, 2021, 2022).
2 Theory and Methods
@@ 2.1 Rigorous Numerics
The natural interval extension of an arithmetical expression is obtained by using interval arithmetic operations instead of exact arithmetic operations at each stage. Different expressions for the same function give rise to different natural interval extensions. These may yield poor approximations due to a phenomenon known as the dependency effect.
Example 1.
If $f(x)=x(1-x)$, then ${\lfloor f\rceil}({[0\!\!1]})={[0\!:\!1]}\times(1-{[0\!:\!1]})={[0\!:\!1]}\times{[0\!:\!1]}={[0\!:\!1]}$. However, $f(x)=x(1-x)=\frac{1}{4}-(\frac{1}{2}-x)^{2}$, and the natural interval extension of this expression yields ${\lfloor f\rceil}({\lfloor 0\!:\!1\rceil})=\frac{1}{4}-(\frac{1}{2}-{[0\!:\!1]})^{2}=\frac{1}{2}-{[-\frac{1}{2}\!:\!+\frac{1}{2}]}^{2}=\frac{1}{4}-{[0\!:\!\frac{1}{4}]}={[0\!:\!\frac{1}{4}]}$, a better approximation.
The dependency effect is particularly strong in the use of Gaussian elimination on intervals to solve linear algebraic equations, which yields very poor results for even moderate-size problems. A viable approach  (Rump, 2010) is to precondition using an approximate inverse $P\approx A^{-1}$ and to use an iterative method such as Gauss-Seidel to solve
@@
A powerful rigorous calculus for continuous functions $\mathbb{R}^{n}\rightarrow\mathbb{R}$ is based around the Taylor models of Makino and Berz (2003). A Taylor model for a function $f:[-1\!:\!+1]^{n}\rightarrow\mathbb{R}$ is a pair $\hat{f}=\langle p,I\rangle$ where $p$ is a polynomial in $n$ variables
@@
Taylor models for smooth functions can be computed using Taylor’s theorem with remainder term. Automatic differentiation (Griewank, 2000) is a technique for computing the derivatives without resorting to symbolic differentiation or divided differences. The basic idea is to evaluate a formula over a data type containing a value $u$ and its (partial) derivatives $\partial{u}/\partial{x}_{j}$ with respect to one or more independent variables. We then use the basic rules of differentiation to propagate the derivatives through the formula. From the basic sum, product and chain rules
@@
To solve the algebraic equation $f(y)=0$, a standard technique is to use the interval Newton operator (Moore, 1966):
@@
There are many methods in the literature for solving differential equations, including (Lohner, 1987; Berz and Makino, 1998; Nedialkov et al., 1999; Zgliczynski, 2002); see also (Tucker, 2011). However, most use some variation on the following procedure to compute the flow $\phi$ of $\dot{x}=f(x)$ starting in an initial set $X_{0}$ at time 00 for a time step $h$. The flow is computable as a fixed-point of the Picard operator
@@
Other problems which have been well-studied in the literature include nonlinear programming (Hansen and Walster, 2004) and constraint propagation (Kearfott, 1996; Jaulin et al., 2001). The latter is used in Ariadne to test emptiness of the intersection of sets.
2.2 Computable Analysis
Computable analysis (Weihrauch, 2000) provides a theoretical foundation for rigorous numerics. An operation is computable if it the result can be rigorously computed to arbitrary accuracy.
@@ 2.2 Computable Analysis
A set may have many admissible respresentations. Two representations are equivalent if it is possible to compute a name for an object in either representation from any name in the other. Sets with equivalence classes of representations we call a (computable) type To emphasise that these types correspond to the standard mathematical spaces, we will sometimes use the terminology mathematical type instead.
Representations of mathematical types are strongly related to topologies on the represented sets. Well-behaved representations of a topological space are quotient maps, and are called admissible quotient representations in (Battenfeld et al., 2007). The topological spaces arising in analysis and geometry all have a unique “natural” representation up to equivalence, so have a canonical computable type. (An “unnatural” representation of the real numbers can be obtained by shifting by an uncomputable constant, but arithmetic is then uncomputable with respect to this representation.) A fundamental result is than only continuous operators can be computable. (Further, we do not know of any “natural” continuous operators which are uncomputable.)
@@
Quotients of countably-based spaces
@@
Since we can construct product and function types, computable types form a Cartesian closed category, which allows constructions of the lambda-calculus; see (Lambek and Scott, 1988).
For a more complete introduction to computable analysis, see (Weihrauch, 2000; Collins, 2020b).
3 Conceptual foundations
@@ 3.2 Types
Many generic classes represent mutually-incompatible sets of objects. For example, Function<Real(RealVector)> class has an argument_size method, and arithmetical operations cannot be applied to scalar functions with different argument sizes. In order to construct a zero function, we need to specify the argument_size; in particular, functions do not have a natural default constructor. The information needed to construct a null element of a class are called characteristics, and CharacteristicsType<Y> is a tuple class containing this data. Numerical classes such as Real have no characteristics.
3.3 Information
@@
An important case of validated information is a finite prefix of a data stream giving effective information. For example, from $\pi=3.14159\cdots$ we can deduce $\pi\in 3.141[59\!:60]$. There is no canonical conversion from effecive to validated information, since we need to specify some accuracy, but we can often perform a binary operation on an effective and a validated object, yielding a validated object. For example, given $\hat{\pi}=3.141[59\!:60]$ and $\epsilon$, we can approximate $e$ to $5$ decimal places, yielding $\hat{\phi}+e=3.141[59\!:60]+2.7182[8\!:9]=5.3134[1\!:\!3]$.
@@ 3.4 Properties
4 Core Functionality: Logic, Numbers, Algebra
@@ 4.1 Logic
4.2 Numbers
In Ariadne, we provide algebraic Integer, Dyadic and Rational classes, based on the GMP library (GMP, 1991-, ). Classes Nat, Nat32, Int and Int32 are wrappers for C++ builtin integral types, and a Decimal class is provided for user input.
@@ 4.2 Numbers
$\begin{gathered}\mathrm{nul}(x)=0,\quad\mathrm{pos}(x)=+x,\quad\mathrm{neg}(x)=-x,\quad\mathrm{sqr}(x)=x^{2},\quad\mathrm{hlf}(x)=x/2,\quad\mathrm{rec}(x)=1/x;\\ \mathrm{add}(x_{1},x_{2})=x_{1}+x_{2},\quad\mathrm{sub}(x_{1},x_{2})=x_{1}-x_{2},\quad\mathrm{mul}(x_{1},x_{2})=x1\times x_{2},\quad\mathrm{div}(x_{1},x_{2})=x_{1}\div x_{2};\\ \mathrm{pow}(x,n)=x^{n},\quad\mathrm{fma}(x_{1},x_{2},y)=(x_{1}\times x_{2})+y;\\[3.0pt] \operatorname{\mathrm{sqrt}}(x),\ \exp(x),\ \log(x),\ \sin(x),\ \cos(x),\ \tan(x),\ \operatorname{\mathrm{atan}}(x);\\ \operatorname{\mathrm{abs}}(x)=|x|,\quad\mathrm{max}(x_{1},x_{2}),\quad\mathrm{min}(x_{1},x_{2});\\ \end{gathered}$
@@
Arithmetic on Bounds<F> and Ball<F,FE> is standard interval arithmetic (see (Moore et al., 2009)), which can easily be implemented in terms of rounded arithmetic.
@@
operator+(Ball<FloatMP,FloatDP>, FloatMP) -> Ball<FloatMP,FloatDP>;
operator-(Real, LowerBound<FloatDP>) -> UpperBound<FloatDP>;
operator*(ValidatedNumber, FloatDP) -> Bounds<FloatDP>;
The templated type alias ArithmeticType<X1,X2> is defined as the result of x1 * x2 - x1 * x2, where x1,x2 have types X1,X2 respectively.
@@ 4.3 Linear algebra
solve(Matrix<X> A, Vector<X> b) -> Matrix<ArithmeticalType<X>>
4.4 Abstract algebra
@@ 5.1 Generic functions
A function with a Scalar arguments is Univariate, and with Vector arguments Multivariate. Hence ValidatedScalarMultivariateFunction is a synonym for the template Function<ValidatedTag,RealScalar(RealVector)>. We also define partially-templated aliases, so ScalarMultivariateFunction<P> is a shorthand for Function<P,RealScalar(RealVector)> and ValidatedFunction<SIG> for Function<ValidatedTag,SIG>.
@@
x = RealVariable("x"); y=RealVariable("y");
f = EffectiveScalarMultivariateFunction([x,y],exp(x)*cos(y));
@@
Ariadne provides extended evaluation operations, so the class ApproximateFunction<Real(Real)> can be evaluated not only on ApproximateNumber, but also on FloatDPApproximation and FloatMPApproximation, and indeed, over any pointwise-algebra over ApproximateNumber. ValidatedFunction<Real(Real)> also supports evaluation on ValidatedNumber, FloatBounds<PR> and FloatBall<PR,PRE>.
@@
5.2 Function patches and models
@@
Figure 1: A scaled function model.
@@
5.3 Polynomial models
The main class of validated function model in Ariadne are the TaylorFunctionModels, based on the Taylor models of Makino and Berz (2003). Since the only existing package implementing Taylor models when Ariadne was under development, COSY Infinity (Makino and Berz, 2006), is not open-source, an implementation within Ariadne itself was developed. Unlike the original Taylor models, where an arbitrary interval $I$ is used as the remainder term, here we only give an error bound $e$, corresponding to an interval remainder $I=[-e\!:\!+e]$.

### v4

@@ Abstract.
Key words and phrases:
Rigorous numerics; Computable analysis; Function calculus; Mathematical software.
1. Introduction
@@
The Ariadne software package (Ariadne , ) provides analysis and verification tools for dynamic systems with continuous state spaces. This includes nonlinear hybrid systems in which continuous evolution is interspersed with discrete events triggered by conditions on the continuous state. This is a very general and complex problem, and requires functionality for many different operations, including solution of algebraic and differential equations and of constrained feasibility problems, and requires general-purpose software. The computational kernel of Ariadne was subsequently developed, based on rigorous numerical methods for working with real numbers, functions and sets in Euclidean space. Previous publications (BBC+, 08; CBGV, 12) have focused on the use of Ariadne for reachability analysis of systems. The purpose of this paper is to describe the low-level functionality of Ariadne that has been used to build up the high-level algorithms, with a focus on the support for working with Euclidean functions. We shall illustrate the functionality with snippets of real C++ code; this code works under Ariadne Release 2.5 (Ariadne, 25).
@@
1.1. Other Tools and Approaches
@@
  • Ariadne (Collins, Geretti, Villa et al., 2002–) for verification of hybrid systems (C++ & Python).
@@
  • iRRAM (Müller, Brauße, 2000–) for real number arithmetic (C++).
@@
  • AERN (Konec̆ný et al., 2005–) for effective real computation (Haskell).
@@
  • ERC (Ziegler, Park et al., 2018–), a language for exact real computation.
The iRRAM package started off as a tool for efficient and straightforward, high-precision computation with real numbers (Mül, 01), but more recent developments have extended the functionality to include a function calculus (BKM, 16) and solution of differential equations. The AERN package is similar in scope to Ariadne, including a function calculus DFKT (14). The ERC language PBC+ (24) has a formal semantics for exact real computation; a motivation for the development of ERC as a computer analysis (as opposed to algebra) system was given in BCZ (22).
@@
  • COSY Infinity (MB, 06), the seminal package in which Taylor models were developed MB (03).
  • CAPD Library (KMWZ, 21), a comprehensive library for analysis of dynamical systems.
  • Flow* (CAS, 13), a tool for hybrid system evolution which also has Taylor models with interval coefficients.
  • JuliaReach (BFF+, 19), a tool for set-based reachability analysis.
  • DynIBEX (AdC, 16), has a custom class for trajectories.
  • CORA (AGK, 18) has Taylor models, but operations use floating-point approximations.
@@
  • Ellipsoidal calculus (KV, 02)
  • Set-valued viability (CQSP, 99)
  • Taylor-ellipsoid models (HVC, 13)
Since the main aim of this paper is to introduce the framework and capabilities of Ariadne, we do not give a full comparison with these other tools here. For a comparison of the performance of Ariadne against other tools on benchmark problems, see (GADSA+, 20, 21, 22).
2. Theory and Methods
@@ 2.1. Rigorous Numerics
The natural interval extension of an arithmetical expression is obtained by using interval arithmetic operations instead of exact arithmetic operations at each stage. Different expressions for the same function give rise to different natural interval extensions. These may yield poor approximations due to a phenomenon known as the dependency effect: For example, if $f(x)=x(1-x)$, then ${\lfloor f\rceil}({[0\!\!1]})={[0\!:\!1]}\times(1-{[0\!:\!1]})={[0\!:\!1]}\times{[0\!:\!1]}={[0\!:\!1]}$. However, $f(x)=x(1-x)=\frac{1}{4}-(\frac{1}{2}-x)^{2}$, and the natural interval extension of this expression yields ${\lfloor f\rceil}({\lfloor 0\!:\!1\rceil})=\frac{1}{4}-(\frac{1}{2}-{[0\!:\!1]})^{2}=\frac{1}{2}-{[-\frac{1}{2}\!:\!+\frac{1}{2}]}^{2}=\frac{1}{4}-{[0\!:\!\frac{1}{4}]}={[0\!:\!\frac{1}{4}]}$, a better approximation.
The dependency effect is particularly strong in the use of Gaussian elimination on intervals to solve linear algebraic equations, which yields very poor results for even moderate-size problems. A viable approach  (Rum, 10) is to precondition using an approximate inverse $P\approx A^{-1}$ and to use an iterative method such as Gauss-Seidel to solve
@@
A powerful rigorous calculus for continuous functions $\mathbb{R}^{n}\rightarrow\mathbb{R}$ is based around the Taylor models of MB (03). A Taylor model for a function $f:[-1\!:\!+1]^{n}\rightarrow\mathbb{R}$ is a pair $\hat{f}=\langle p,I\rangle$ where $p$ is a polynomial in $n$ variables
@@
Taylor models for smooth functions can be computed using Taylor’s theorem with remainder term. Automatic differentiation (Gri, 00) is a technique for computing the derivatives without resorting to symbolic differentiation or divided differences. The basic idea is to evaluate a formula over a data type containing a value $u$ and its (partial) derivatives $\partial{u}/\partial{x}_{j}$ with respect to one or more independent variables. We then use the basic rules of differentiation to propagate the derivatives through the formula. From the basic sum, product and chain rules
@@
To solve the algebraic equation $f(y)=0$, a standard technique is to use the interval Newton operator (Moo, 66):
@@
There are many methods in the literature for solving differential equations, including (Loh, 87; BM, 98; NJC, 99; Zgl, 02); see also (Tuc, 11). However, most use some variation on the following procedure to compute the flow $\phi$ of $\dot{x}=f(x)$ starting in an initial set $X_{0}$ at time 00 for a time step $h$. The flow is computable as a fixed-point of the Picard operator
@@
Other problems which have been well-studied in the literature include nonlinear programming (HW, 04) and constraint propagation (Kea, 96; JKDW, 01). The latter is used in Ariadne to test emptiness of the intersection of sets.
2.2. Computable Analysis
Computable analysis (Wei, 00) provides a theoretical foundation for rigorous numerics. An operation is computable if it the result can be rigorously computed to arbitrary accuracy.
@@ 2.2. Computable Analysis
A set may have many admissible respresentations. Two representations are equivalent if it is possible to compute a name for an object in either representation from any name in the other. Sets with equivalence classes of admissible representations we call a (computable) type To emphasise that these types correspond to the standard mathematical spaces, we will sometimes use the terminology mathematical type instead.
Representations of mathematical types are strongly related to topologies on the represented sets. Well-behaved representations of a topological space are quotient maps, and are called admissible quotient representations in (BSS, 07). The topological spaces arising in analysis and geometry all have a unique “natural” representation up to equivalence, so have a canonical computable type. (An “unnatural” representation of the real numbers can be obtained by shifting by an uncomputable constant, but arithmetic is then uncomputable with respect to this representation.) A fundamental result is than only continuous operators can be computable. (Further, we do not know of any “natural” continuous operators which are uncomputable.)
@@
Quotients of countably-based spaces:
@@
Since we can construct product and function types, computable types form a Cartesian closed category, which allows constructions of the lambda-calculus; see (LS, 88).
For a more complete introduction to computable analysis, see (Wei, 00; Col20a, ).
3. Conceptual foundations
@@ 3.2. Types
Many classes represent mutually-incompatible sets of objects. For example, the class Vector<Real> has a size() method, and arithmetic cannot be applied to vectors with different sizes. In order to construct a zero vector, we need to specify the size; in particular, vectors do not have a natural default constructor. The information needed to construct a null element of a class are called characteristics, and CharacteristicsType<Y> is a tuple class containing this data. Numerical classes such as Real have no characteristics.
3.3. Information
@@
An important case of validated information is a finite prefix of a data stream giving effective information. For example, from $\pi=3.14159\cdots$ we can deduce $\pi\in 3.141[59\!:60]$. There is no canonical conversion from effecive to validated information, since we need to specify some accuracy, but we can often perform a binary operation on an effective and a validated object, yielding a validated object. For example, given $\hat{\pi}=3.141[59\!:60]$ and $\mathrm{e}$, we can approximate $\mathrm{e}$ to $5$ decimal places, yielding $\hat{\pi}+\mathrm{e}=3.141[59\!:60]+2.7182[8\!:9]=5.3134[1\!:\!3]$.
@@ 3.4. Properties
4. Core Functionality: Logic, Numbers, Algebra
@@ 4.1. Logic
4.2. Numbers
In Ariadne, we provide algebraic Integer, Dyadic and Rational classes, based on the GMP library (GMP , ). Classes Nat, Nat32, Int and Int32 are wrappers for C++ builtin integral types, and a Decimal class is provided for user input.
@@ 4.2. Numbers
$\begin{gathered}\mathrm{nul}(x)=0,\quad\mathrm{pos}(x)=+x,\quad\mathrm{neg}(x)=-x,\quad\mathrm{sqr}(x)=x^{2},\quad\mathrm{hlf}(x)=x/2,\quad\mathrm{rec}(x)=1/x;\\ \mathrm{add}(x_{1},x_{2})=x_{1}\!+\!x_{2},\ \ \mathrm{sub}(x_{1},x_{2})=x_{1}\!-\!x_{2},\ \ \mathrm{mul}(x_{1},x_{2})=x1\!\times\!x_{2},\ \ \mathrm{div}(x_{1},x_{2})=x_{1}\!\div\!x_{2};\\ \mathrm{pow}(x,n)=x^{n},\quad\mathrm{fma}(x_{1},x_{2},y)=(x_{1}\times x_{2})+y;\\[3.0pt] \operatorname{\mathrm{sqrt}}(x),\ \exp(x),\ \log(x),\ \sin(x),\ \cos(x),\ \tan(x),\ \operatorname{\mathrm{atan}}(x);\\ \operatorname{\mathrm{abs}}(x)=|x|,\quad\mathrm{max}(x_{1},x_{2}),\quad\mathrm{min}(x_{1},x_{2});\\ \end{gathered}$
@@
Arithmetic on Bounds<F> and Ball<F,FE> is standard interval arithmetic (see (MKC, 09)), which can easily be implemented in terms of rounded arithmetic.
@@
operator+(Ball<F,FE>, F) -> Ball<F,FE>;
operator-(Real, LowerBound<F>) -> UpperBound<F>;
operator*(ValidatedNumber, Float<PR>) -> Bounds<Float<PR>>;
The templated type alias ArithmeticType<X1,X2> is defined as the result of the expression x1 * x2 - x1 * x2, where x1,x2 have types X1,X2 respectively.
@@ 4.3. Linear algebra
solve(Matrix<X> A, Vector<X> b)
-> Vector<ArithmeticalType<X>>
4.4. Abstract algebra
@@ 5.1. Generic functions
A function with a Scalar argument is Univariate, and with a Vector argument Multivariate. Hence ValidatedScalarMultivariateFunction is a synonym for the template Function<ValidatedTag,RealScalar(RealVector)>. We also define partially-templated aliases, so ValidatedFunction<SIG> is short for Function<ValidatedTag,SIG> and ScalarMultivariateFunction<P> for Function<P,RealScalar(RealVector)>.
@@
x = RealVariable("x"); t=RealVariable("t");
f = EffectiveScalarMultivariateFunction([x,t],x*exp(t));
@@
Extended evaluation operations are provided, so ApproximateFunction<Real(Real)> can be evaluated not only on ApproximateNumber, but also on FloatDPApproximation and FloatMPApproximation, and indeed, over any pointwise-algebra over ApproximateNumber. Similarly, ValidatedFunction<Real(Real)> also supports evaluation on ValidatedNumber, FloatBounds<PR> and FloatBall<PR,PRE>.
@@
5.2. Function patches and models
@@
Figure 1. A scaled function model.
@@
5.3. Polynomial models
The main class of validated function model in Ariadne are the TaylorFunctionModels, based on the Taylor models of MB (03). Since the only existing package implementing Taylor models when Ariadne was under development, COSY Infinity (MB, 06), is not open-source, an implementation within Ariadne itself was developed. Unlike the original Taylor models, where an arbitrary interval $I$ is used as the remainder term, here we only give an error bound $e$, corresponding to an interval remainder $I=[-e\!:\!+e]$.